OpenStudy (anonymous):

Find the length of the longer leg of a 30°-60°-90° triangle with a hypotenuse of 6 in.

OpenStudy (anonymous):

The short leg is opposite the smaller angle, so Sin30 = short/hypotenuse => hypotenuse = short/(sin30) Sin30 = 1/2, so hyp = short/(1/2) = short * 2